OpenStudy (anonymous):

NO ONE IS HELPING TODAY AT ALL I ASKED LIKE A THOUSAND TIMES AND NOTHING PLZ HELP

OpenStudy (anonymous):

A car travels 1 6 of the distance between two cities in 3 5 of an hour. At this rate, what fraction of the distance between the two cities can the car travel in 1 hour

OpenStudy (anonymous):

The students at Midtown Middle school sold flowers as a fundraiser in September and October. In October, they charged $1.50 for each flower. The October price was a 20% increase of the September price. Part A What was the price of the flowers in September

OpenStudy (anonymous):

who paid most and by how much

OpenStudy (anonymous):

plz help me no one has

OpenStudy (anonymous):

People are helping :/

OpenStudy (anonymous):

srry my internet mess up

OpenStudy (anonymous):

there wernt before i open dis question

OpenStudy (anonymous):

ik i got 1.20 but idk who pay more

OpenStudy (anonymous):

o wait a sec

OpenStudy (anonymous):

The seventh-grade class earned 40% of the selling price of each flower. In September, they sold 900 flowers. In October, they sold 700 flowers.

OpenStudy (anonymous):

i need to know who paid more and by how much i cant get it

OpenStudy (anonymous):

i tried everything but none of them are right idk if it the number i put 1.20 or wat i cant seem to get it right and then there another question i need help with plz help me on those to

OpenStudy (kropot72):

The car travels 1/6 of the distance in 3/5 of an hour, therefore the distance traveled in 1/5 of an hour is: \[\large \frac{\frac{1}{6}}{3}\] In one hour the fraction of distance traveled is given by: \[\large \frac{1}{6}\times\frac{5}{3}=you\ can\ calculate\]

OpenStudy (anonymous):

ok hang on

OpenStudy (anonymous):

ok thx i need help with one more plz come back